OpenStudy (anonymous):

before the clowns arrived, only 35% of the children had happy faces. If 117 children did not have happy faces, how many children were there in all?

OpenStudy (anonymous):

180

OpenStudy (anonymous):

how would i write that in a ratio box? @eHoaX

OpenStudy (anonymous):

so if 35% of them are happy then the sad one would be 65% because the max percent is 100%. so we know 117 are sad let x represent the total amount of kids x*0.65=117 solve for x and you get 180

OpenStudy (anonymous):

@eHoaX how would i set the corresponding proportion, and set the equation up algebraically, and solve it?

OpenStudy (anonymous):

so 65% corresponds to 117 and 35% corresponds to an unknown number so \[\frac{ 35 }{ 65 }=\frac{ x }{ 117 }\] where x is the number of happy kids

OpenStudy (anonymous):

this x is different from the other x in order to get the total amount you have to add it to 117